OpenStudy (anonymous):

solve this system of equations using the elemination method: 3x-2y=2 4x-7y=33

OpenStudy (anonymous):

please show work

OpenStudy (sasogeek):

3x-2y = 2 .... (lets call this (e1) ) 4x-7y = 33 .... (lets call this (e2) ) we need to eliminate one of the variables in order to find the other... elimination method... but we can't do that if the variable has different coefficients in both equations... so we can play with the numbers so that the coefficients evolve into becoming the same number... MAGIC! :P here's what i mean lets say we want to eliminate the "x" variable... we have to make the coefficients of the "x" in both equations the same... here's how... 4(3x-2y) = 4(2) .......multiply the whole of equation 1 by 4 (lets call this (e3) ) 3(4x-7y) = 3(33) .......multiply the whole of equation 2 by 3 (lets call this (e4) ) from (e3) 12x-8y = 8 ......... (e5) from (e4) 12x-21y = 99 ....... (e6) eliminating x (subtracting (e6) from (e5) ) 12x - 8y = 8 12x - 21y = 99 0 +13y = -91 13y = -91 y = -91/13 y = -7 if y=-7 .... can you find x on your own?

OpenStudy (anonymous):

no

OpenStudy (karatechopper):

so u mean he went thru all tht for nothing????????????????????

OpenStudy (karatechopper):

"MAGIC!" lol

OpenStudy (pokemon23):

awesome ;)

OpenStudy (anonymous):

Nice work saso

OpenStudy (sasogeek):

substitute the value of y (which is -7) into one of the equations and solve for x

OpenStudy (anonymous):

still lost sorry

OpenStudy (sasogeek):

where ever u see y in the first equation, put (-7), then solve for x

OpenStudy (anonymous):

i'm still confused about the first equation

OpenStudy (sasogeek):

take it easy... relax, sit back and read through the solution i posted first, over and over... 2 times... then tell me if you're still lost :)

OpenStudy (anonymous):

i keep getting the answer: -6.2

OpenStudy (sasogeek):

ok ummm 3x - 2y = 2 but y == -7 so 3x - 2(-7) = 2 3x - (-14) = 2 3x + 14 = 2 3x = 2-14 3x = -12 x = -12/3 x = -4

OpenStudy (sasogeek):

how did u solve it to get -6.2? i wanna know what u did... i might be able to tell where u went wrong...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3x-2y=2 4x-7y=33 i did 4 times 3.. and made the top 12 negative -12x-2=2 12x-7y=33 then for the top i did -12 time -2, and -12 times 2. i did the same thing for the bottom line, except different numbers. 24=-24 -84=396 ( the twelves canceled out) then i had 24 and -84, and -24 with 396 and i got: -60=372 then i divided those 2 numbers. which gives me: -6.2 this is the way may teacher taught me. please help

OpenStudy (sasogeek):

o.o... so wrong... take a look at my explanation a couple of times... you'll get ur head around it, i don't even understand ur explanation but i'll look at it for a while, maybe i'll know what u were up to

OpenStudy (karatechopper):

lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4(3x-2y) = 4(2) .......multiply the whole of equation 1 by 4 (lets call this (e3) ) 3(4x-7y) = 3(33) .......multiply the whole of equation 2 by 3 (lets call this (e4) ) i don't understand this part

OpenStudy (anonymous):

cus you explain this part to me?

OpenStudy (anonymous):

can not cus